OpenStudy (anonymous):

Find the measure of each interior angle of a regular convex dodecagon. 100(10-2)=200;40° 360(12-2)=3600;300° 90(12-2)=900;75° 1800(12-2)=1800;150°

OpenStudy (beastboy17):

are there options or is this it

OpenStudy (anonymous):

That's it

OpenStudy (beastboy17):

okay so do you have a suspicion of what the answer could be

OpenStudy (welshfella):

the sum of the exterior angles = 360 1 exterior angle = 360 / 12 = 30 degrees so interior angle 180 - 30

OpenStudy (anonymous):

A. 100(10-2)=200;40° B. 360(12-2)=3600;300° C. 90(12-2)=900;75° D. 1800(12-2)=1800;150°

OpenStudy (anonymous):

I think it could be D

OpenStudy (beastboy17):

id probably agree but im not 100%

OpenStudy (beastboy17):

id hate to agree and have it be incorrect

OpenStudy (welshfella):

formula for sum of all angles = 180(n - 2) here n = 12 dividing this by 12 will give the measure of one angle

OpenStudy (beastboy17):

@Michele_Laino

OpenStudy (welshfella):

1808 10 = 1800 1800 / 12 = ?

OpenStudy (welshfella):

* first line should read 180 * 10

OpenStudy (welshfella):

what 's 1800 divided by 12?